OpenStudy (anonymous):

Find the arc length of the major arc. Use 3.14 @aum @M.D.Saxon

OpenStudy (anonymous):

(Let Me Attach The File First)

OpenStudy (m.d.saxon):

ok

OpenStudy (anonymous):

Here's the pic that goes to this question

OpenStudy (anonymous):

@phi can u help me with my question when u get the time

OpenStudy (anonymous):

@aum can u help me with my question when u get the time

OpenStudy (aum):

Major arc is 360 - 85 = 275 degrees. The circumference of a circle is 2*pi*r and the arc is 360 degrees. That is, 360 degrees corresponds to an arc length of 2*pi*r. Therefore, 275 degrees corresponds to an arc length of 275 / 360 * 2*pi*r (by proportions). Put r = 13.

OpenStudy (anonymous):

The correct answer must also be rounded to two decimal places

OpenStudy (aum):

When you plug in the numbers into your calculator what do you get?

OpenStudy (anonymous):

i got 62.39552076

OpenStudy (aum):

round it two decimal places.

OpenStudy (anonymous):

62.39 or 62.40

OpenStudy (aum):

Yes, 62.40.

OpenStudy (anonymous):

OK!!! THANX @aum

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

oh no @aum 62.40 isn't the right answer my program said "Try Again"

OpenStudy (aum):

That is what I am getting. The 85 degrees is the minor arc and (360-85 = 275) degrees is the major arc.

OpenStudy (aum):

Okay, I think I used the calculator pi which is to more decimal places and here they are asking you to use 3.14 for pi. If I do that I get arc length = 62.3638 which when rounded to 2 decimal places gives 62.36. Try 62.36

OpenStudy (anonymous):

Ok thank you @aum